# Pentagram maps over rings, Grassmannians, and skewers

### v1

@@ 1 Introduction
The pentagram map has been generalized in several different directions. One can intersect diagonals joining vertices that are further apart, consider polygons in higher-dimensional spaces, etc. [10, 11, 6]. Here, we are interested in different kinds of generalizations, which are based on replacing conventional notions of points and lines by more general objects with similar properties. One example of such a map is the Grassmann pentagram map, defined in [4] and further studied in [13, 2]. In this setting, the role of points is played by elements of $\mathrm{Gr}(k,3k)$, i.e. $k$-dimensional subspaces in a $3k$-dimensional vector space, while lines are elements of $\mathrm{Gr}(2k,3k)$. Any two generic points $V_{1},V_{2}\in\mathrm{Gr}(k,3k)$ belong to a single line $V_{1}\oplus V_{2}\in\mathrm{Gr}(2k,3k)$, and any two generic lines $W_{1},W_{2}\in\mathrm{Gr}(2k,3k)$ meet at a point $W_{1}\cap W_{2}\in\mathrm{Gr}(k,3k)$. The Grassmann pentagram map takes a (cyclically labeled) sequence of points $V_{i}\in\mathrm{Gr}(k,3k)$, which we think of as a polygon is the Grassmannian $\mathrm{Gr}(k,3k)$, to a new sequence given by $(V_{i}\oplus V_{i+2})\cap(V_{i+1}\oplus V_{i+3})$. In other words, just like the classical pentagram map, the Grassmann map is defined by intersecting diagonals connecting second-nearest vertices. Furthermore, the classical pentagram map is nothing else but the $k=1$ version of the Grassmann map.
@@
Furthermore, we prove that the pentagram map is integrable for any stably finite ground ring $R$, thus generalizing the results of [4, 13] on integrability of the Grassmann map and answering the question of [20] regarding integrability of the skewer map. Integrability is understood in this paper as the existence of a Lax form. A discrete dynamical system $\phi$ on a set $X$ (i.e., a map $\phi\colon X\to X$) is said to admit a Lax form if to every $x\in X$ one can assign an operator in such a way that operators assigned to $x$ and $\phi(x)$ are similar. In the setting of pentagram maps over a ring $R$, these operators are pseudo-difference operators with coefficients in $R$.
@@
In Section 2, we recall the definition and standard properties of stably finite rings. Details can be found, e.g., in [12, Chapter 1, §1B].
Section 3 serves as an introduction into projective geometry over rings. We note that while ring geometry is a well-studied subject (see [23]), the setting of stably finite rings appears to be new. In Section 3.1, we define projective spaces as projectivizations of arbitrary modules. The case of free modules is treated in detail in Section 3.3. We call the projectivization of a free rank $d+1$ module the projective space of dimension $d$. Sections 3.2 and 3.4 are devoted to projective spaces over the full matrix ring, which is the first of our two main examples. The second main example, the ring of dual numbers, is discussed in Section 3.5.
@@
In Section 5, we introduce our main technical tool: difference operators with coefficients in a ring. After discussing basic properties of such operators, we use that language to give an algebraic description of the moduli space of polygons in the projective plane over a ring. The main result of the section, Theorem 5.7, is a ring counterpart of [8, Proposition 3.3] which gives a description of real polygons in terms of difference operators. We note that the proof in the ring case is substantially more technical due to the lack of commutativity and essentially relies on the stable finiteness condition.
@@
Acknowledgments. This work was supported by NSF grant DMS-2008021.
@@ 2 Stably finite rings
In this section, we recall the definition and basic properties of stably finite rings – the main class of rings relevant to the present paper; for details see [12, Chapter 1, §1B]. In what follows, all rings are assumed to be unital and non-trivial (i.e. have a multiplicative identity element $1\neq 0$), but may be non-commutative and/or have zero divisors. All modules are supposed to be unitary.
@@
    the ring $\mathrm{End}_{R}(M)$ is Dedekind finite (recall that a ring is Dedekind finite if $ab=1$ implies $ba=1$).
@@
$2\Rightarrow 3$. Suppose the ring $\mathrm{End}_{R}(M)$ is Dedekind finite, and let $\phi\colon M\to M\oplus N$ be an isomorphism. Let also $\pi\colon M\oplus N\to M$ be the projection onto the first summand. Then, since $\pi$ is surjective, and $i\colon M\to M\oplus N$ be the inclusion map. Then $(\pi\phi)(\phi^{-1}i)=\pi i=\mathrm{id}$, so, by Dedekind finiteness, $\pi\phi$ is an isomorphism, and $N=\mathrm{Ker}\,\pi=0$.
@@
Suppose $R$ is a ring, and $M$ is a free $R$-module of finite rank. Then $M$ is Hopfian if and only if its dual module $M^{*}=\mathrm{Hom}_{R}(M,R)$ is Hopfian.
@@
For a free module of finite rank, the map sending an endomorphism to its adjoint is an anti-isomorphism between the rings $\mathrm{End}_{R}(M)$ and $\mathrm{End}_{R}(M^{*})$. Since the Dedekind finiteness condition is left-right symmetric, it is preserved by anti-isomorphisms. ∎
Furthermore, if $M$ is free of finite rank (i.e., admits a finite basis), then the Hopfian condition can be restated as follows.
@@
This is equivalent to saying that $R^{n}\simeq R^{m}$ (as left or right modules) implies $m=n$. Assume that $\phi\colon R^{n}\to R^{m}$ is an isomorphism. Without loss of generality, let $n\leq m$. Then the projection map $\pi\colon R^{m}\to R^{n}$, $\pi(r_{1},\dots,r_{m}):=(r_{1},\dots,r_{n})$ is surjective, and so is the composition $\pi\phi\colon R^{n}\to R^{n}$. Therefore, $\pi\phi$ is an isomorphism, and $\pi$ is injective. So, $m=n$. ∎
@@ 3.1 The projectivization of a module
Let $M$ be a (left or right) $R$-module. A subspace $N\subset M$ is a free submodule which admits a direct complement in $M$ (i.e. there is a submodule $K\subset M$, not necessarily free, such that $M=N\oplus K$).
@@
Let $M$ be an $R$-module. Then rank one subspaces of $M$ are precisely cyclic submodules of $M$ generated by unimodular elements.
@@
We give a proof for the case of left modules, the case of right modules is analogous. Suppose $P\subset M$ is a rank one subspace. Then, by definition, $P$ is freely generated over $R$ by some $v\in M$, and is a direct summand in $M$, i.e. $M=P\oplus N$. Therefore, for each $w\in M$ there exists unique $r\in R$ and $u\in N$ such that $w=rv+u$. Define a homomorphism $\xi\colon M\to R$ by $\xi(rv+u):=r$. Then $\xi(v)=1$, ensuring $v$ is unimodular.
@@
Let $M$ be an $R$-module. Two unimodular elements $v,w\in M$ generate the same subspace if and only if they are unit multiples of each other.
@@
We give a proof for the case of left modules, the case of right modules is analogous. Clearly, if $r$ is a unit, then $Rv=Rrv$. Conversely, assume that $Rv=Rw$. Then $w\in Rv$, so $w=rv$ for some $r\in R$. Likewise, $v=r^{\prime}w$ for some $r^{\prime}\in R$. So, $w=rv=rr^{\prime}w$. Since $w$ is unimodular, this implies $rr^{\prime}=1$. Similarly, $r^{\prime}r=1$. So, $r$ is a unit. ∎
@@
Let $M$ be an $R$-module. There is a one-to-one correspondence between points of $\mathbb{P}(M)$ and unimodular elements in $M$ modulo units.
@@ 3.2 Projective spaces over the full matrix ring
Here we consider our first main example – the projective space over the full matrix ring, cf. [23, Section 7] and [21].
Let $\mathbb{F}$ be a field, and $R=\text{M}_{k}(\mathbb{F})$ be the ring of $k\times k$ matrices over $\mathbb{F}$. By definition, projective spaces over $R$ are projectivizations of $R$-modules. Since $R$ is isomorphic to its opposite ring via transposition, there is no distinction between left and right modules. So, we will restrict our attention to left modules. Any left module $M$ over $\text{M}_{k}(\mathbb{F})$ is isomorphic to the one of the form $V^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}V$, where $V$ is a (not necessarily finite-dimensional) vector space over $\mathbb{F}$, and $\text{M}_{k}(\mathbb{F})$ acts on the first tensor factor. The correspondence $V\mapsto V^{k}$ provides additive equivalence between the categories of $\mathbb{F}$-modules (vector spaces) and $\text{M}_{k}(\mathbb{F})$-modules. This is also known as the Morita equivalence between the corresponding rings $\mathbb{F}$ and $\text{M}_{k}(\mathbb{F})$. Thus, for the purpose of describing projective spaces over $\text{M}_{k}(\mathbb{F})$, it suffices to consider left modules of the form $M=V^{k}$, where $V$ is a vector space over $\mathbb{F}$.
@@
Submodules of the module $V^{k}$ are of the form $W^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}W$ where $W\subset V$ is an arbitrary subspace of $V$. The correspondence $W\mapsto W^{k}$ gives an isomorphism between the lattice of subspaces of $V$ and the lattice of submodules of $V^{k}$. In particular, any submodule of $V^{k}$ admits a direct complement, which means that subspaces of $V^{k}$ are precisely its free submodules. Further, $W^{k}$ is a free rank one $\text{M}_{k}(\mathbb{F})$-module precisely when $\dim_{\mathbb{F}}W=k$, so rank one free submodules of $V^{k}$ are precisely of the form $W^{k}$ for $W\in\mathrm{Gr}(k,V)$. ∎
@@ 3.4 The projective plane over the full matrix ring
In words, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ can be described as follows: points are $k$-dimensional subspaces of $\mathbb{F}^{3k}$, lines are $2k$-dimensional subspaces of $\mathbb{F}^{3k}$, while incidence of points and lines corresponds to containment. Note that two generic points $P_{1},P_{2}\in\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ lie on a unique line $P_{1}\oplus P_{2}$. Likewise, two generic lines $L_{1},L_{2}$ meet at a unique point $L_{1}\cap L_{2}$. However, if $\dim P_{1}\oplus P_{2}<2k$, then there exist infinitely many lines containing points $P_{1},P_{2}$ (and, if $\dim L_{1}\cap L_{2}>k$, there exist infinitely many points lying on $L_{1},L_{2}$). Thus, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ is not a projective plane in the sense of incidence geometry. Incidence spaces with properties similar to those of $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ are known as Barbilian planes, see [3, 22]. It is shown in [22] that such planes are precisely the projective planes over rings of stable rank two, of which the matrix ring is an example.
@@ 3.5 The projective plane over the dual numbers
Let $\mathcal{A}(V)$ be the space of affine straight lines in $V$. We define an incidence structure on $\mathcal{A}(V)$ as follows: both points and lines are $\mathcal{P}=\mathcal{L}=\mathcal{A}(V)$ and a point is incident to a line if and only if the affine lines intersect and are perpendicular to each other. In this geometry, both the operation of joining two points with a line and intersecting two lines boil down to taking the common perpendicular, or skewer, of the corresponding affine lines.
We will show that there is natural isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$. The corresponding mapping $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ between the point spaces is a variant of the Study mapping, see [15, Section 2.3] and [18]. For an affine line $\{p+tv\mid t\in\mathbb{R}\}$, where $p,v\in V$ and $v\neq 0$, let
@@
In particular, for any $v\neq 0$ we have that $\zeta(v,p)$ is unimodular and therefore $\mathbb{D}\zeta(v,p)$ is a subspace. Moreover, this subspace is independent on the choice of the direction vector $v$ and base point $p$ of the line $v+tp$, since for any $\lambda\in\mathbb{R}$ we have $\zeta(\lambda v,p)=\lambda\zeta(v,p)$ and $\zeta(v,p+\lambda v)=\zeta(v,p)$. So, the mapping $\phi_{P}$ is well-defined.
@@
A $\mathbb{D}$-module is an $\mathbb{R}[\varepsilon]$-module annihilated by $\varepsilon^{2}$. Therefore, any finitely generated $\mathbb{D}$-module $M$ is isomorphic to a unique module of the form $M^{p,q}=\mathbb{D}^{\oplus p}\oplus(\mathbb{D}/(\varepsilon))^{\oplus q}$ (this is just the Jordan decomposition over the reals for the multiplication by $\varepsilon$ in $M$). Such a module is free if and only if $q=0$. So, the direct sum of two modules, at least one of which is not free, cannot be free. ∎
@@
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Let $N\subset M$ be a subspace of rank $k$. Then $N^{\bot}:=\{v\in M\mid\langle v,w\rangle=0\mbox{ for all }w\in N\}$ is also a subspace of rank $n-k$.
@@
Consider the restriction mapping $\rho\colon M^{*}\to N^{*}$. Since $N$ is free of rank $k$, so is $N^{*}$, and $M^{*}=\mathrm{Ker}\,\rho\oplus K$ for some submodule $K\subset M^{*}$. Thus, by Proposition 3.17, the submodule $\mathrm{Ker}\,\rho\subset M$ is free. Furthermore, $\rho_{K}\colon K\to N^{*}$ is an isomorphism, so $K$ has rank $k$ and $\mathrm{Ker}\,\rho$ has rank $n-k$. Now, consider the mapping $\phi\colon M\to M^{*}$ given by the inner product. Since the inner product is non-degenerate, the mapping $\phi$ is injective. Therefore, since $M$ is a finite-dimensional real algebra, $\phi$ is, in fact, an isomorphism. Together, $N^{\bot}=\phi^{-1}(\mathrm{Ker}\,\rho)$ is a free submodule of $M$ which has rank $n-k$ and admits a direct complement, $\phi^{-1}(K)$, as needed. ∎

### v2

@@ 1 Introduction
The pentagram map has been generalized in several different directions. One can intersect diagonals joining vertices that are further apart, consider polygons in higher-dimensional spaces, etc. [10, 11, 6]. Here, we are interested in other kinds of generalizations, which are based on replacing conventional notions of points and lines by more general objects with similar properties. One example of such a map is the Grassmann pentagram map, defined in [4] and further studied in [14, 2]. In this setting, the role of points is played by elements of $\mathrm{Gr}(k,3k)$, i.e. $k$-dimensional subspaces in a $3k$-dimensional vector space, while lines are elements of $\mathrm{Gr}(2k,3k)$. Any two generic points $V_{1},V_{2}\in\mathrm{Gr}(k,3k)$ belong to a single line $V_{1}\oplus V_{2}\in\mathrm{Gr}(2k,3k)$, and any two generic lines $W_{1},W_{2}\in\mathrm{Gr}(2k,3k)$ meet at a point $W_{1}\cap W_{2}\in\mathrm{Gr}(k,3k)$. The Grassmann pentagram map takes a (cyclically labeled) sequence of points $V_{i}\in\mathrm{Gr}(k,3k)$, which we think of as a polygon in the Grassmannian $\mathrm{Gr}(k,3k)$, to a new sequence given by $(V_{i}\oplus V_{i+2})\cap(V_{i+1}\oplus V_{i+3})$. In other words, just like the classical pentagram map, the Grassmann map is defined by intersecting diagonals connecting second-nearest vertices. Furthermore, the classical pentagram map is nothing else but the $k=1$ version of the Grassmann map.
@@
Furthermore, we prove that the pentagram map is integrable for any stably finite ground ring $R$, thus generalizing the results of [4, 14] on integrability of the Grassmann map and answering the question of [21] regarding integrability of the skewer map. Integrability is understood in this paper as the existence of a Lax form. A discrete dynamical system $\phi$ on a set $X$ (i.e., a map $\phi\colon X\to X$) is said to admit a Lax form if to every $x\in X$ one can assign an operator in such a way that the operators assigned to $x$ and $\phi(x)$ are similar. In the setting of pentagram maps over a ring $R$, these operators are pseudo-difference operators with coefficients in $R$. When $R$ is a field, we recover the Lax form of the classical pentagram map given in [6, 8].
@@
In Section 2, we recall the definition and standard properties of stably finite rings. Details can be found, e.g., in [13, Chapter 1, §1B].
Section 3 serves as an introduction into projective geometry over rings. We note that while ring geometry is a well-studied subject (see [24]), the setting of stably finite rings appears to be new. In Section 3.1, we define projective spaces as projectivizations of arbitrary modules. The case of free modules is treated in detail in Section 3.3. We call the projectivization of a free rank $d+1$ module the projective space of dimension $d$. We note that these projective spaces are not projective spaces in the sense of incidence geometry (also known as combinatorial projective spaces). In particular, for a general (stably finite) ring $R$, there are pairs of points in $\mathbb{P}^{2}(R)$ which are not contained in a unique line, as well as lines that do not share a unique point.
Sections 3.2 and 3.4 are devoted to projective spaces over the full matrix ring, which is the first of our two main examples. The second main example, the ring of dual numbers, is discussed in Section 3.5.
@@
In Section 5, we introduce our main technical tool: difference operators with coefficients in a ring. After discussing basic properties of such operators, we use that language to give an algebraic description of the moduli space of polygons in the projective space over a ring. The main result of the section, Theorem 5.7, is a ring counterpart of [8, Proposition 3.3] which gives a description of real polygons in terms of difference operators. We note that the proof in the ring case is substantially more technical due to the lack of commutativity and essentially relies on the stable finiteness condition.
@@
Acknowledgments. The authors are grateful to Boris Khesin, Nicholas Ovenhouse, Richard Schwartz, and Max Weinreich for fruitful discussions. This work was supported by NSF grant DMS-2008021.
@@ 2 Stably finite rings
In this section, we recall the definition and basic properties of stably finite rings – the main class of rings relevant to the present paper; for details see [13, Chapter 1, §1B]. In what follows, all rings are assumed to be unital and non-trivial (i.e. have a multiplicative identity element $1\neq 0$), but may be non-commutative and/or have zero divisors. All modules are supposed to be unitary.
@@
    the ring $\mathrm{End}_{R}(M)$ is Dedekind-finite (recall that a ring is Dedekind-finite if $ab=1$ implies $ba=1$).
@@
$2\Rightarrow 3$. Suppose the ring $\mathrm{End}_{R}(M)$ is Dedekind-finite, and let $\phi\colon M\to M\oplus N$ be an isomorphism. Let also $\pi\colon M\oplus N\to M$ be the projection onto the first summand, and $i\colon M\to M\oplus N$ be the inclusion map. Then $(\pi\phi)(\phi^{-1}i)=\pi i=\mathrm{id}$, so, by Dedekind-finiteness, $\pi\phi$ is an isomorphism, and $N=\mathrm{Ker}\,\pi=0$.
@@
Suppose $R$ is a ring, and $M$ is a free $R$-module of finite rank (i.e., admits a finite basis). Then $M$ is Hopfian if and only if its dual module $M^{*}=\mathrm{Hom}_{R}(M,R)$ is Hopfian.
@@
For a free module of finite rank, the map sending an endomorphism to its adjoint is an anti-isomorphism between the rings $\mathrm{End}_{R}(M)$ and $\mathrm{End}_{R}(M^{*})$. Since the Dedekind-finiteness condition is left-right symmetric, it is preserved by anti-isomorphisms. ∎
Furthermore, if $M$ is free of finite rank, the Hopfian condition can be restated as follows.
@@
This is equivalent to saying that $R^{m}\simeq R^{n}$ (as left or right modules) implies $m=n$. Assume that $\phi\colon R^{n}\to R^{m}$ is an isomorphism. Without loss of generality, let $n\leq m$. Then the projection map $\pi\colon R^{m}\to R^{n}$, $\pi(r_{1},\dots,r_{m}):=(r_{1},\dots,r_{n})$ is surjective, and so is the composition $\pi\phi\colon R^{n}\to R^{n}$. Therefore, $\pi\phi$ is an isomorphism, and $\pi$ is injective. So, $m=n$. ∎
@@ 3.1 The projectivization of a module
Let $M$ be a (left or right) $R$-module. A subspace $N\subset M$ is a free submodule which admits a direct complement in $M$ (i.e., there is a submodule $K\subset M$, not necessarily free, such that $M=N\oplus K$).
@@
Rank one subspaces of an $R$-module $M$ are precisely cyclic submodules of $M$ generated by unimodular elements.
@@
We give a proof for the case of left modules (the case of right modules is analogous). Suppose $P\subset M$ is a rank one subspace. Then, by definition, $P$ is freely generated over $R$ by some $v\in M$, and is a direct summand in $M$, i.e. $M=P\oplus N$. Therefore, for each $w\in M$ there exists unique $r\in R$ and $u\in N$ such that $w=rv+u$. Define a homomorphism $\xi\colon M\to R$ by $\xi(rv+u):=r$. Then $\xi(v)=1$, ensuring $v$ is unimodular.
@@
Two unimodular elements $v,w$ of an $R$-module $M$ generate the same subspace if and only if they are unit multiples of each other.
@@
We give a proof for the case of left modules (the case of right modules is analogous). Clearly, if $r$ is a unit, then $Rv=Rrv$. Conversely, assume that $Rv=Rw$. Then $w\in Rv$, so $w=rv$ for some $r\in R$. Likewise, $v=r^{\prime}w$ for some $r^{\prime}\in R$. So, $w=rv=rr^{\prime}w$. Since $w$ is unimodular, this implies $rr^{\prime}=1$. Similarly, $r^{\prime}r=1$. So, $r$ is a unit. ∎
@@
For any $R$-module $M$, there is a one-to-one correspondence between points of $\mathbb{P}(M)$ and unimodular elements in $M$ modulo units.
@@ 3.2 Projective spaces over the full matrix ring
Here we consider our first main example – the projective space over the full matrix ring, cf. [24, Section 7] and [22].
Let $\mathbb{F}$ be a field, and $R=\text{M}_{k}(\mathbb{F})$ be the ring of $k\times k$ matrices over $\mathbb{F}$. By definition, projective spaces over $R$ are projectivizations of $R$-modules. Since $R$ is isomorphic to its opposite ring via transposition, there is no distinction between left and right modules. So, we will restrict our attention to left modules. Any left module $M$ over $\text{M}_{k}(\mathbb{F})$ is isomorphic to the one of the form $V^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}V$, where $V$ is a (not necessarily finite-dimensional) vector space over $\mathbb{F}$, and $\text{M}_{k}(\mathbb{F})$ acts on the first tensor factor. The correspondence $V\mapsto V^{k}$ provides an additive equivalence between the categories of $\mathbb{F}$-modules (vector spaces) and $\text{M}_{k}(\mathbb{F})$-modules. This is also known as the Morita equivalence between the corresponding rings $\mathbb{F}$ and $\text{M}_{k}(\mathbb{F})$. Thus, for the purpose of describing projective spaces over $\text{M}_{k}(\mathbb{F})$, it suffices to consider left modules of the form $M=V^{k}$, where $V$ is a vector space over $\mathbb{F}$.
@@
Submodules of the module $V^{k}$ are of the form $W^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}W$ where $W\subset V$ is an arbitrary subspace of $V$. The correspondence $W\mapsto W^{k}$ gives an isomorphism between the lattice of subspaces of $V$ and the lattice of submodules of $V^{k}$. In particular, any submodule of $V^{k}$ admits a direct complement, which means that subspaces of $V^{k}$ are precisely its free submodules. Further, $W^{k}$ is a free rank one $\text{M}_{k}(\mathbb{F})$-module precisely when $\dim_{\mathbb{F}}W=k$, so rank one free submodules of $V^{k}$ are those of the form $W^{k}$ for $W\in\mathrm{Gr}(k,V)$. ∎
@@ 3.4 The projective plane over the full matrix ring
In words, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ can be described as follows: points are $k$-dimensional subspaces of $\mathbb{F}^{3k}$, lines are $2k$-dimensional subspaces of $\mathbb{F}^{3k}$, while incidence of points and lines corresponds to containment. Note that two generic points $P_{1},P_{2}\in\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ lie on a unique line $P_{1}\oplus P_{2}$. Likewise, two generic lines $L_{1},L_{2}$ meet at a unique point $L_{1}\cap L_{2}$. However, if $\dim P_{1}\oplus P_{2}<2k$, then there exist infinitely many lines containing points $P_{1},P_{2}$ (and, if $\dim L_{1}\cap L_{2}>k$, there exist infinitely many points lying on $L_{1},L_{2}$). Thus, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ is not a combinatorial projective plane. Incidence spaces with properties similar to those of $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ are known as Barbilian planes, see [3, 23]. It is shown in [23] that such planes are precisely the projective planes over rings of stable rank two, of which the matrix ring is an example.
@@ 3.5 The projective plane over the dual numbers
Let $\mathcal{A}(V)$ be the space of affine straight lines in $V$. We define an incidence structure on $\mathcal{A}(V)$ as follows: both points and lines are affine lines in $V$, i.e., $\mathcal{P}=\mathcal{L}=\mathcal{A}(V)$, and a point is incident to a line if and only if the affine lines intersect and are perpendicular to each other. In this geometry, both the operation of joining two points with a line and intersecting two lines boil down to taking the common perpendicular, or skewer, of the corresponding affine lines.
We will show that there is natural isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$. The corresponding mapping $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ between the point spaces is a variant of the Study mapping, see [16, Section 2.3] and [19]. For an affine line $\{p+tv\mid t\in\mathbb{R}\}$, where $p,v\in V$ and $v\neq 0$, let
@@
In particular, for any $v\neq 0$ we have that $\zeta(v,p)$ is unimodular and therefore $\mathbb{D}\zeta(v,p)$ is a subspace. Moreover, this subspace does not depend on the choice of the direction vector $v$ and base point $p$ of the line $v+tp$, since for any $\lambda\in\mathbb{R}$ we have $\zeta(\lambda v,p)=\lambda\zeta(v,p)$ and $\zeta(v,p+\lambda v)=\zeta(v,p)$. So, the mapping $\phi_{P}$ is well-defined.
@@
A $\mathbb{D}$-module is an $\mathbb{R}[\varepsilon]$-module annihilated by $\varepsilon^{2}$. Therefore, any finitely generated $\mathbb{D}$-module $M$ is isomorphic to a unique module of the form $M^{p,q}=\mathbb{D}^{\oplus p}\oplus(\mathbb{D}/(\varepsilon))^{\oplus q}$ (this is just the Jordan decomposition over the reals for the multiplication by $\varepsilon$ in $M$). Such a module is free if and only if $q=0$. So, the direct sum of two finitely generated $\mathbb{D}$-modules, at least one of which is not free, cannot be free. ∎
@@
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Let $N\subset M$ be a subspace of rank $k$. Then $N^{\bot}:=\{v\in M\mid\langle v,w\rangle=0\mbox{ for all }w\in N\}$ is also a subspace, and $\mathrm{rank}\,N^{\bot}=n-k$.
@@
Consider the restriction mapping $\rho\colon M^{*}\to N^{*}$. Since $N$ is free of rank $k$, so is $N^{*}$, and $M^{*}=\mathrm{Ker}\,\rho\oplus K$ for some submodule $K\subset M^{*}$. Thus, by Proposition 3.17, the submodule $\mathrm{Ker}\,\rho\subset M{{}^{*}}$ is free. Furthermore, $\rho_{K}\colon K\to N^{*}$ is an isomorphism, so $K$ has rank $k$ and $\mathrm{Ker}\,\rho$ has rank $n-k$. Now, consider the mapping $\phi\colon M\to M^{*}$ given by the inner product. Since the inner product is non-degenerate, the mapping $\phi$ is injective. Therefore, since $M$ is a finite-dimensional real algebra, $\phi$ is, in fact, an isomorphism. Together, $N^{\bot}=\phi^{-1}(\mathrm{Ker}\,\rho)$ is a free submodule of $M$ which has rank $n-k$ and admits a direct complement, $\phi^{-1}(K)$, as needed. ∎